# KVComm: Enabling Efficient LLM Communication through Selective KV Sharing

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly deployed in multi-agent systems, where effective inter-model communication is crucial. Existing communication protocols either rely on natural language, incurring high inference costs and information loss, or on hidden states, which suffer from information concentration bias and inefficiency. To address these limitations, we propose KVComm, a novel communication framework that enables efficient communication between LLMs through selective sharing of KV pairs. KVComm leverages the rich information encoded in the KV pairs while avoiding the pitfalls of hidden states. We introduce a KV layer-wise selection strategy based on attention importance scores with a Gaussian prior to identify the most informative KV pairs for communication. Extensive experiments across diverse tasks and model pairs demonstrate that KVComm achieves comparable performance to the upper-bound method, which directly merges inputs to one model without any communication, while transmitting as few as 30% of layers’ KV pairs. Our study highlights the potential of KV pairs as an effective medium for inter-LLM communication, paving the way for scalable and efficient multi-agent systems.
@@ 1 Introduction
However, natural language communication leads to high inference costs due to the need for multiple decoding steps, and may not fully capture the rich information that needs to be shared between LLMs as information is lost in the sampling process (Pham et al., 2023; Ramesh & Li, 2025) that occurs as each new token is produced. To address this limitation, recent works have explored alternative communication protocols that leverage the internal representations of LLMs. CIPHER (Pham et al., 2023) proposed to use the embedding space as the medium of communication between LLMs. Namely, they pass the weighted average of the token embeddings from one LLM to another, facilitating more efficient information exchange. Rather than using the embedding space, AC (Ramesh & Li, 2025) transmits the intermediate activations, specifically the last token’s hidden state. They replace the last token’s hidden state of the receiver’s model ($\mathcal{M}_{r}$) with that of the sender’s model ($\mathcal{M}_{s}$), allowing a more direct transfer of information. While these methods have shown promising results, they still face challenges in terms of communication efficiency and effectiveness. CIPHER (Pham et al., 2023) still requires multiple decoding steps, which can be costly, and AC (Ramesh & Li, 2025) may lead to information loss as only limited activation information is transmitted.
We start with the question: What is the most effective way to communicate between LLMs? We argue that an ideal communication protocol should satisfy the following criteria: ① Effectiveness: It should enable $\mathcal{M}_{r}$ to effectively utilize the information from $\mathcal{M}_{s}$. ② Efficiency: It should minimize the computation needed by $\mathcal{M}_{s}$ and the amount of data transmitted between models. ③ Generality: It should be applicable to a wide range of tasks and model architectures, ensuring its versatility in different scenarios. We choose to use activation information as the medium of communication, as no decoding steps are needed for $\mathcal{M}_{s}$, and $\mathcal{M}_{r}$ can directly utilize the rich information encoded in the activations. We study different types of activation information (i.e., hidden states and KV pairs), and in Section 2.2, we show that hidden states suffer from information concentration bias, where the last token’s hidden state contains most of the information needed for the model’s output. This makes it challenging to design an effective communication protocol using the last token’s hidden state. Furthermore, we find that using all tokens’ hidden states does not guarantee effective communication. A dilemma arises: if the hidden states are taken from the early layers of $\mathcal{M}_{s}$, the computation benefit is limited since the computation cost is similar to concatenating the two inputs; if the hidden states are prepended to the later layers of $\mathcal{M}_{r}$, the performance drops significantly.
@@ 2.1 Problem Formulation
We formally define the problem of solving a contextual task through the communication of two LLMs: $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$. $\mathcal{M}_{s}$ takes as input a context $C$, and generates the required information $I_{C}$ to be communicated. $\mathcal{M}_{r}$ takes as input the query $Q$ and the information $I_{C}$ from $\mathcal{M}_{s}$, and produces the final output. In this work, we limit the choices of the two LLMs to (1) two instances of the same LLM, and (2) two models that are fine-tuned versions of the same base LLM. The goal is to design an efficient communication protocol that allows $\mathcal{M}_{s}$ to effectively convey the necessary information to $\mathcal{M}_{r}$ while minimizing the amount of data transmitted.
@@ 2.2.1 Token Importance at Different Positions
To ensure efficient communication with hidden states built on this observation, two conditions must hold: (1) $\mathcal{M}_{s}$ must send at least the last token’s hidden state, and (2) the communication protocol should preserve $\mathcal{M}_{r}$’s last token state as much as possible. The protocol in Ramesh & Li (2025) either replaces $\mathcal{M}_{r}$’s last token state with that of $\mathcal{M}_{s}$ or averages the two, but both cause information loss in $\mathcal{M}_{r}$’s last token state, harming its performance.
@@
Figure 3: Prepending hidden states is not effective unless hidden states are from and to the early layers.
@@ 3.2 KV Selection Strategies
Another hypothesis H2 is that KV pairs from layers exhibiting stronger attention distributions are more effective for communication. Intuitively, if a head consistently allocates high attention mass to the given tokens, its KV cache encodes salient contextual relations that are critical for the model’s reasoning. Attention concentration thus serves as a proxy for the communication value of a KV subset, suggesting that such heads should be prioritized for selection. This hypothesis is also validated by our experiments in Section 4.5.
@@
$\hat{S}_{a}^{l}=\frac{1}{HT}\sum_{h=1}^{H}\sum_{t=1}^{T}\sum_{c=1}^{C}a_{h,t,c}^{l},$
where $H$ is the number of attention heads, $T$ is the number of tokens in the query, $C$ is the number of context tokens, and $a_{h,t,c}^{l}$ is the attention weight assigned by head $h$ at layer $l$ from token $t$ to context token $c$. $\hat{S}_{a}^{l}$ is then normalized to the range $[0,1]$ across all layers to obtain the final attention importance score $S_{a}^{l}=\frac{\hat{S}_{a}^{l}-\min_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}}{\max_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}-\min_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}}$.
@@
For each model pair and dataset, the top $M$ layers are selected based on the selection scores computed from a calibration set. The selected layers are then fixed and used for all samples in the test set. We found that a calibration set as small as a single sample is sufficient to obtain a robust selection that generalizes well to the entire test set, as shown in the experiments in Appendix G.
@@ 4.1 Experimental Setup
We evaluate KVComm on a diverse set of contextual reasoning tasks. Following Ramesh & Li (2025), we synthetically generate two datasets, Countries, which asks questions about countries based on landmark information, and Tipsheets, which requires investment decisions from financial tips. Examples of these two datasets are shown in Table 3 in Section B.1. Moreover, we select six benchmarks, including HotpotQA (Yang et al., 2018), QASPER (Dasigi et al., 2021), MuSiQuest (Trivedi et al., 2022), two subsets of LongBench (Bai et al., 2024)(MultiFieldQA-en and 2WikiMQA), and TMATH (Qi et al., 2025). The last dataset is a mathematical problem-solving dataset that contains hints as context. We use ROUGE-L Recall as the evaluation metric for the last dataset, and F1 score for all other datasets. Statistics are summarized in Table 4 in Section B.1.
@@
We compare KVComm with several representative approaches: Baseline (no communication between $\mathcal{M}_{r}$ and $\mathcal{M}_{s}$), Skyline (concatenating context $C$ and query $Q$ as an upper bound), Natural Language Debate (NLD) (Du et al., 2023), CIPHER (Pham et al., 2023), and AC (Ramesh & Li, 2025). Detailed descriptions for these methods are provided in Section B.4. Implementation details are provided in Section B.2.
@@ 4.2 Communication Results
Table 1 reports results on three model pairs fine-tuned from the same base LLM. The results on other model pairs are provided in LABEL:tab:more_communication_results in Appendix E, which show similar trends. We observe that KVComm consistently outperforms all baseline communication methods across datasets and model pairs. AC can outperform the Baseline method on some datasets, but they are still significantly worse than KVComm and Skyline, as hidden states of $\mathcal{M}_{r}$ are corrupted during communication. NLD and CIPHER methods perform poorly on most datasets, as they are designed for improving the answer quality through debate rather than communication. $\mathcal{M}_{s}$ cannot effectively convey useful information to $\mathcal{M}_{r}$ through debate, as $\mathcal{M}_{s}$ has no information about the query $Q$. Our KVComm framework can achieve comparable performance than the Skyline method when selecting 70% of layers’ KV pairs for communication, demonstrating the effectiveness of our selection strategy. Even when selecting only 30% of layers’ KV pairs, KVComm can still outperform most baseline communication methods on many datasets, showing its potential for efficient communication with minimal overhead.
@@ 4.3 Benefit of Selective KV Over One Contiguous Chunk
Figure 6 shows that using a single contiguous chunk for communication yields good performance only in a small region of the hyperparameter space, making it tricky to find the right hyperparameters. In contrast, the scatter and curve plots in Figure 6 demonstrate that KVComm consistently achieves the best or even outperforms the best contiguous chunk setting for the same number of layers. Line plots in Figure 6 show that contiguous chunks are most effective when taken from intermediate layers, consistent with hypothesis H1 in Section 3.2. All results are on HotpotQA with the Llama-3.1-8B pair, with more in Appendix I.
@@ 4.4 Ablation Study on Selection Strategy
Table 2 compares KVComm with random selection. We find that KVComm consistently outperforms random selection across different datasets and selection ratios. When the ratio is high (i.e., 0.7), the performance gap between our selection strategy and random selection becomes smaller, as more layers are selected and the impact of the selection strategy is reduced. However, when the ratio is low (i.e., 0.3), our selection strategy significantly outperforms random selection, demonstrating its effectiveness in selecting the most informative layers for communication. Comparison results on other model pairs are provided in LABEL:tab:more_comparison_with_random in Appendix F, which show similar trends.
@@ 4.6 System Efficiency
Figure 8: KVComm requires less computation compared to Skyline.
@@ 5 Related Work
Lots of work has focused on accelerating LLM inference. Computation-level methods such as FlashAttention (Dao et al., 2022) and Memory-Efficient Attention (Rabe & Staats, 2021) reduce memory and speed up attention; system-level methods such as vLLM (Kwon et al., 2023) and DeepSpeed-Inference (Aminabadi et al., 2022) improve overall throughput and latency; and model-level methods such as quantization (Lin et al., 2024) and pruning (Ma et al., 2023) reduce model size and complexity. These approaches are orthogonal to ours and can be combined with KVComm to further improve efficiency.
@@
DroidSpeak (Liu et al., 2024b) also explores KV cache reuse across LLMs, but with key differences. Their goal is to accelerate inference for multiple queries with the same prefix, while we focus on efficient communication between two LLMs on contextual tasks. They reuse a single contiguous chunk of layers and recompute the rest, whereas our strategy flexibly selects non-contiguous layers based on attention importance and a Gaussian prior. Moreover, their method incurs extra overhead from recomputation, while ours directly integrates the selected KV pairs into $\mathcal{M}_{r}$ without recomputation. Despite the different problem settings, we compare their contiguous-chunk strategy with ours in Section 4.3, showing the advantages of our approach.
@@
CIPHER (Pham et al., 2023) proposed using embedding space as the medium of communication. They pass the weighted average of the token embeddings from one LLM to another. Moreover, AC (Ramesh & Li, 2025) proposed to use the last token’s hidden state as the medium of communication. They replace the last token’s hidden state of the receiver model with that of the sender model. Instead, we propose to use the KV pairs as the medium, which can preserve more information than just using the last token’s hidden state. We also propose a more effective selection strategy for choosing which KV pairs to share, which can further improve efficiency.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly deployed in multi-agent systems, where effective inter-model communication is crucial. Existing communication protocols either rely on natural language, incurring high inference costs and information loss, or on hidden states, which suffer from information concentration bias and inefficiency. To address these limitations, we propose KVComm, a novel communication framework that enables efficient communication between LLMs through selective sharing of KV pairs. KVComm leverages the rich information encoded in the KV pairs while avoiding the pitfalls of hidden states. We introduce a KV layer-wise selection strategy based on attention importance scores with a Gaussian prior to identify the most informative KV pairs for communication. Extensive experiments across diverse tasks and model pairs demonstrate that KVComm achieves comparable performance to the upper-bound method, which directly merges inputs to one model without any communication, while transmitting as few as 30% of layers’ KV pairs. Our study highlights the potential of KV pairs as an effective medium for inter-LLM communication, paving the way for scalable and efficient multi-agent systems [^1] Our code and data can be found at https://github.com/Zephyroam/KVComm.
@@ 1 Introduction
However, natural language communication leads to high inference costs due to the need for multiple decoding steps, and may not fully capture the rich information that needs to be shared between Large Language Models, as information is lost in the sampling process (Pham et al., 2023; Ramesh and Li, 2025) that occurs as each new token is produced. To address this limitation, recent works have explored alternative communication protocols that leverage the internal representations of LLMs. CIPHER (Pham et al., 2023) proposed to use the embedding space as the medium of communication between LLMs. Namely, they pass the weighted average of the token embeddings from one LLM to another, facilitating more efficient information exchange. Rather than using the embedding space, AC (Ramesh and Li, 2025) transmits the intermediate activations, specifically the last token’s hidden state. They replace the last token’s hidden state of the receiver’s model ($\mathcal{M}_{r}$) with that of the sender’s model ($\mathcal{M}_{s}$), allowing a more direct transfer of information. While these methods have shown promising results, they still face challenges in terms of communication efficiency and effectiveness. CIPHER (Pham et al., 2023) still requires multiple decoding steps, which can be costly, and AC (Ramesh and Li, 2025) may lead to information loss as only limited activation information is transmitted.
We start with the question: What is the most effective way to communicate between LLMs? We argue that an ideal communication protocol should satisfy the following criteria: ① Effectiveness: It should enable $\mathcal{M}_{r}$ to effectively utilize the information from $\mathcal{M}_{s}$. ② Efficiency: It should minimize the computation needed by $\mathcal{M}_{s}$ and the amount of data transmitted between models. ③ Generality: It should be applicable to a wide range of tasks and model architectures, ensuring its versatility in different scenarios. We choose to use activation information as the medium of communication, as no decoding steps are needed for $\mathcal{M}_{s}$, and $\mathcal{M}_{r}$ can directly utilize the rich information encoded in the activations. We study different types of activation information (i.e., hidden states and KV pairs), and in Section 2.2, we show that hidden states suffer from information concentration bias, where the last token’s hidden state contains most of the information needed for the model’s output. This makes it challenging to design an effective communication protocol using the last token’s hidden state. Furthermore, we find that using all tokens’ hidden states from a single layer of the sender $\mathcal{M}_{s}$ does not guarantee effective communication. A dilemma arises: if the hidden states are taken from the early layers of $\mathcal{M}_{s}$, the computation benefit is limited since the computation cost is similar to concatenating the two inputs; if the hidden states are prepended to the later layers of $\mathcal{M}_{r}$, the performance drops significantly.
@@ 2.1 Problem Formulation
We formally define the problem of solving a contextual task through the communication of two LLMs: $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$. $\mathcal{M}_{s}$ takes as input a context $C$, and generates the required information $I_{C}$ to be communicated. $\mathcal{M}_{r}$ takes as input the query $Q$ and the information $I_{C}$ from $\mathcal{M}_{s}$, and produces the final output. In this work, we limit the choices of the two LLMs to (1) two instances of the same LLM, and (2) two models that are fine-tuned versions of the same base LLM. The objective is to design a communication protocol that jointly optimizes the communication, computation efficiency, and the information fidelity between $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$.
@@ 2.2.1 Token Importance at Different Positions
To ensure efficient communication with hidden states built on this observation, two conditions must hold: (1) $\mathcal{M}_{s}$ must send at least the last token’s hidden state, and (2) the communication protocol should preserve $\mathcal{M}_{r}$’s last token state as much as possible. The protocol in Ramesh and Li (2025) either replaces $\mathcal{M}_{r}$’s last token state with that of $\mathcal{M}_{s}$ or averages the two, but both cause information loss in $\mathcal{M}_{r}$’s last token state, harming its performance.
@@
Figure 3: Prepending helps only when using early-layer hidden states.
@@ 3.2 KV Selection Strategies
Another hypothesis H2 is that KV pairs from layers exhibiting stronger attention distributions are more effective for communication. We quantify this notion of attention distribution using the attention importance score $S_{a}^{l}$, defined in Equation 1 below. We deem layer $l_{i}$ to exhibit stronger attention distribution than $l_{j}$, if $S_{a}^{l_{i}}>S_{a}^{l_{j}}$. Intuitively, if a head consistently allocates high attention mass to the given tokens, its KV cache encodes salient contextual relations that are critical for the model’s reasoning. Attention concentration thus serves as a proxy for the communication value of a KV subset, suggesting that such heads should be prioritized for selection. This hypothesis is also validated by our experiments in Section 4.5.
@@
$\hat{S}_{a}^{l}=\frac{1}{HT}\sum_{h=1}^{H}\sum_{t=1}^{T}\sum_{c=1}^{|C|}a_{h,t,c}^{l},$ (1)
where $H$ is the number of attention heads, $T$ is the number of tokens in the query, $|C|$ is the number of context tokens, and $a_{h,t,c}^{l}$ is the attention weight assigned by head $h$ at layer $l$ from token $t$ to context token $c$. $\hat{S}_{a}^{l}$ is then normalized to the range $[0,1]$ across all layers to obtain the final attention importance score $S_{a}^{l}=\frac{\hat{S}_{a}^{l}-\min_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}}{\max_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}-\min_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}}$.
@@
For each model pair and dataset, the top $M$ layers are selected based on the selection scores computed from a calibration set. The selected layers are then fixed and used for all samples in the test set. We found that a calibration set as small as a single sample is sufficient to obtain a robust selection that generalizes well to the entire test set, as shown in the experiments in LABEL:appendix:calibration_set_size.
@@ 4.1 Experimental Setup
We evaluate KVComm on a diverse set of contextual reasoning tasks. Following Ramesh and Li (2025), we synthetically generate two datasets, Countries, which asks questions about countries based on landmark information, and Tipsheets, which requires investment decisions from financial tips. Examples of these two datasets are shown in Table 3 in Section B.1. Moreover, we select six benchmarks, including HotpotQA (Yang et al., 2018), QASPER (Dasigi et al., 2021), MuSiQuest (Trivedi et al., 2022), two subsets of LongBench (Bai et al., 2024)(MultiFieldQA-en and 2WikiMQA), and TMATH (Qi et al., 2025). The last dataset is a mathematical problem-solving dataset that contains hints as context. We use ROUGE-L Recall as the evaluation metric for the last dataset, and F1 score for all other datasets. Statistics are summarized in Table 4 in Section B.1.
@@
We compare KVComm with several representative approaches: Baseline (no communication between $\mathcal{M}_{r}$ and $\mathcal{M}_{s}$), Skyline (concatenating context $C$ and query $Q$ as an upper bound), Natural Language Debate (NLD) (Du et al., 2023), CIPHER (Pham et al., 2023), and AC (Ramesh and Li, 2025). Detailed descriptions for these methods are provided in Section B.4. Implementation details are provided in Section B.2.
@@ 4.2 Communication Results
Table 1 reports results on three model pairs fine-tuned from the same base LLM. The results on other model pairs are provided in LABEL:tab:more_communication_results in Appendix F, which show similar trends. We observe that KVComm consistently outperforms all baseline communication methods across datasets and model pairs. AC can outperform the Baseline method on some datasets, but they are still significantly worse than KVComm and Skyline, as hidden states of $\mathcal{M}_{r}$ are corrupted during communication.
NLD and CIPHER can achieve performance close to that of KVComm or Skyline on Countries and Tipsheets datasets, which is because these datasets require only a very small and highly salient amount of information to be transferred. For all other datasets, the sender has access to the entire context but not the question, and natural-language communication cannot reliably extract and transmit the task-relevant subset of information. As a result, NLD and CIPHER perform substantially below KVComm on complex, long-context reasoning tasks. We conduct further experiments in LABEL:appendix:communication_length to eliminate the influence of hyperparameters.
KVComm can achieve comparable performance to Skyline when selecting 70% of layers’ KV pairs for communication, demonstrating the effectiveness of our selection strategy. Even when selecting only 30% of layers’ KV pairs, KVComm can still outperform most baseline communication methods on many datasets, showing its potential for efficient communication with minimal overhead.
@@ 4.3 Benefit of Selective KV Over One Contiguous Chunk
Figure 6 shows that using a single contiguous chunk for communication yields good performance only in a small region of the hyperparameter space, making it tricky to find the right hyperparameters. In contrast, the scatter and curve plots in Figure 6 demonstrate that KVComm consistently achieves the best or even outperforms the best contiguous chunk setting for the same number of layers. Line plots in Figure 6 show that contiguous chunks are most effective when taken from intermediate layers, consistent with hypothesis H1 in Section 3.2. All results are on HotpotQA with the Llama-3.1-8B pair, with more in LABEL:appendix:one_chunk.
@@ 4.4 Ablation Study on Selection Strategy
Table 2 compares KVComm with random selection. We find that KVComm consistently outperforms random selection across different datasets and selection ratios. When the ratio is high (i.e., 0.7), the performance gap between our selection strategy and random selection becomes smaller, as more layers are selected and the impact of the selection strategy is reduced. However, when the ratio is low (i.e., 0.3), our selection strategy significantly outperforms random selection, demonstrating its effectiveness in selecting the most informative layers for communication. Comparison results on other model pairs are provided in LABEL:tab:more_comparison_with_random in LABEL:appendix:more_comparison_with_random, which show similar trends.
@@ 4.6 System Efficiency
In addition to FLOPs, we also report the memory consumption among methods. KVComm similarly shows a substantial memory advantage over Skyline, as the reduced number of communicated layers not only lowers computation but also alleviates memory pressure. On Tipsheets, KVComm uses 23% to 73% less memory than Skyline. KVComm This further highlights the efficiency of our KVComm framework in achieving lightweight inter-model communication.
@@
Figure 8: KVComm requires less computation and memory compared to Skyline.
@@ 5 Related Work
Lots of work has focused on accelerating LLM inference. Computation-level methods such as FlashAttention (Dao et al., 2022) and Memory-Efficient Attention (Rabe and Staats, 2021) reduce memory and speed up attention; system-level methods such as vLLM (Kwon et al., 2023) and DeepSpeed-Inference (Aminabadi et al., 2022) improve overall throughput and latency; and model-level methods such as quantization (Lin et al., 2024) and pruning (Ma et al., 2023) reduce model size and complexity. These works mainly focus on working with only one model processing a single long input with the aim of minimizing computation cost. These approaches are orthogonal to ours and can be combined with KVComm to further improve efficiency.
@@
DroidSpeak (Liu et al., 2024b) aims to accelerate inference for queries with shared prefixes. It reuses the partial KV cache of these prefixes among different queries. Specifically, it empirically selects a single contiguous chunk of layers and recomputes the rest with large calibration overhead, whereas our strategy flexibly selects non-contiguous layers with low overhead, without needing to recompute the remaining layers. Despite different problem settings, we compare their contiguous-chunk strategy with ours in Section 4.3, showing the advantages of our approach.
Ye et al. (2025) adjusts KV cache for shared content by referencing a pool of cached examples-termed anchors that store observed cache deviations under varying prefixes. Our work goes beyond this related work by: 1) enabling a different type of communication, where the receiver does not have access to the context, 2) making it possible to efficiently and selectively choose layers of KV pairs that will be transmitted, and 3) being able to work effectively across different models that are fine-tuned from one model.
@@
CIPHER (Pham et al., 2023) proposed using embedding space as the medium of communication. They pass the weighted average of the token embeddings from one LLM to another. Moreover, AC (Ramesh and Li, 2025) proposed to use the last token’s hidden state as the medium of communication. They replace the last token’s hidden state of the receiver model with that of the sender model. Instead, we propose to use the KV pairs as the medium, which can preserve more information than just using the last token’s hidden state. We also propose a more effective selection strategy for choosing which KV pairs to share, which can further improve efficiency.
@@
6 Discussion
In this section, we discuss the limitations, clarify the scope of current design choices, and outline promising directions for future research. Additional discussions can be found in LABEL:appendix:more_discussion.
Heterogeneous Model Architectures
Our current KVComm framework assumes that both LLMs share the same base architecture, i.e., identical models or fine-tuned versions of the same base LLM. This is because KV pair structures differ substantially across model families, making direct KV exchange undefined. This architecture dependency is a practical limitation but not a fundamental one. Future work could explore learning latent projections, adapters, or other transformation functions to enable KV exchange across heterogeneous architectures.
Multiple Sender/Receiver Extensions
While we focus on a single sender-receiver pair in this work, KVComm can be naturally extended to multiple senders and/or receivers. KVComm can integrate information from multiple senders by concatenating KV caches, and multiple receivers can independently select layers based on their own attention patterns. As shown in LABEL:appendix:multi_agent, we mathematically extend our framework to multiple senders, and perform a preliminary experiment with two senders and one receiver, showing that multiple senders can improve performance due to diversified information sources. However, a systematic study of scaling behaviors in larger multi-agent networks remains future work.
Context-adaptive Online Calibration
KVComm currently adopts a fixed layer-selection strategy after calibration for simplicity and computational efficiency, while context-adaptive selection is a promising extension. KVComm can naturally support online and dynamic selection. A demonstration and analysis of this mechanism is provided in LABEL:appendix:online_calibration.
Layer Selection Priors
Given our goal of keeping the method simple, efficient, and broadly reproducible, we opt for the Gaussian prior. Other alternatives, such as entropy-weighted or data-driven, are promising but introduce significantly higher complexity, e.g., larger calibration sets, training a selector, or risking overfitting to a particular task distribution. Exploring more sophisticated priors is an interesting direction for future work.